OpenStudy (babynini):

u = 3i+2j v=5i-j a) find the angle between u and v b) find the component of u along v c) find proj_v u

OpenStudy (babynini):

@jim_thompson5910 I did part a but it came out wrong o.o

OpenStudy (babynini):

Angle: cos(theta)= (u*v)/|u||v|

OpenStudy (babynini):

(u*v) = 17 right?

jimthompson5910 (jim_thompson5910):

u dot v should be 13

jimthompson5910 (jim_thompson5910):

you lost a negative sign

OpenStudy (babynini):

hrm. Yeah I added 2 instead of subtracting it.

OpenStudy (babynini):

|u| = sq13 ?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

how about |v|

OpenStudy (babynini):

|v| = sq26

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (babynini):

so then cos(theta) = 13/(13 sq2)

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (babynini):

theta = 45

jimthompson5910 (jim_thompson5910):

good

OpenStudy (babynini):

ah k, it was just that 13

OpenStudy (babynini):

component of u along v would be 2.54?

jimthompson5910 (jim_thompson5910):

sorry got distracted

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

I'm getting 2.5495097567964

jimthompson5910 (jim_thompson5910):

did you get about the same?

OpenStudy (babynini):

mhm!

OpenStudy (babynini):

Sorry, I didn't see you had replied.

jimthompson5910 (jim_thompson5910):

that's ok I didn't see it earlier too

jimthompson5910 (jim_thompson5910):

did you figure out part c)?

OpenStudy (babynini):

Yep I got 5/2i - 1/2 j

jimthompson5910 (jim_thompson5910):

I got <2.5, -0.5> which is the exact same

OpenStudy (babynini):

Fantastic :) the form you wrote it in is probably better haha

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (babynini):

oki! thanks :)